# Neural Isometries: Taming Transformations for Equivariant ML

### v1

@@ 1 Introduction
In this paper, we propose Neural Isometries (NIso), an autoencoder framework which learns to map the observation space to a latent space in which encodings are related by tractable, highly-structured linear maps whenever their corresponding observations are geometrically related in world space.
Specifically, observations are encoded into latent functions by preserving their spatial dimensions. Images, for instance, can be encoded into latents defined over a lower resolution grid or the patch tokens of a ViT. For observations sharing some potentially unknown relationship in world space, we enforce their encodings be related by a functional map $\tau$ – a linear transformation on the space of latent functions. In particular, we require that $\tau{}$ is an isometry such that it preserves a learned inner product and commutes with a learned functional operator, in the sense that rigid body transformations commute with the Laplace operator in Euclidean space as do Lorentz transformations with the d’Alembert operator in Minkowski space.
Neural Isometries form an architecture-agnostic regime for self-supervised equivariant representation learning. To this end, we experimentally validate two principle claims regarding the efficacy and general applicability of our approach:
@@ 2 Related Work
Essential to our approach is the parameterization of transformations between observations in the latent space not as group representations, but instead as functional maps. Introduced in the seminal work of Ovsjanikov et al. [18], functional maps (FMs) provide a powerful medium for interpreting, manipulating, and constructing otherwise intractable mappings between 3D shapes via their realization as linear transformations on spaces of functions on meshes, forming the basis for state-of-the-art pipelines in shape matching and correspondence [19, 20, 21, 22]. Beyond 3D shapes, FMs can be seen as a tool to parameterize transformations between observations viewed as functions, in the sense that images, for instance, are functions that map 2D pixel coordinates to RGB colors. Integral to their study and implementation is the Laplace-Beltrami operator (the generalization of the Laplace operator to function spaces on manifolds), and in particular the expression of FMs in its eigenbasis which can expose specific geometric properties of deformations. In particular, isometries (distance-preserving transformations) manifest as highly-structured matrices not unlike IRs, being orthogonal and commuting with the diagonal matrix of eigenvalues and are thus approximately block-diagonal. That said, FMs are recovered through regularized linear solves and as such lack the immutability inherent in the analytical expressions for IRs of compact groups. However, freed from nagging theoretical constraints, approximately isometric FMs and their orthogonal relaxations have displayed remarkable representational capacity, well modeling a variety of highly-complex non-rigid deformations [23] including those without group structure, such as partial [24] and inter-genus correspondence [25]. Please see [26] for an outstanding introduction to functional maps.
@@
A number of recent approaches have proposed autoencoder frameworks wherein given symmetries in the base space manifest as simple operations in the latent space [27, 28, 29, 30, 31, 32, 33]. Perhaps most similar to our approach is recent work on a Neural Fourier Transform (NFT) that has sought to manifest group actions in observation space as IRs in the the latent space  [34, 35]. These methods offer impressive theoretical guarantees under the conditions that the observed transformations are either known or are a group action, although these assumptions may not hold for real-world data with complex symmetries. In contrast, our method is wholly unsupervised, and assumes no knowledge of the transformations in observation space nor that they even form a group. Furthermore, in lieu of an extensive theoretical justification we instead seek to validate the efficacy of our approach experimentally, including in geometry processing and real-world 3D vision tasks.
Figure 2: Overview of Neural Isometries (NIso). NIso learn a latent space where transformations of observations manifest as isometries, achieved by regularizing the functional maps $\tau$ between latents to commute with a learned operator $\Omega$, parameterized via its spectral decomposition into a mass matrix $\mathbb{M}$, eigenfunctions $\Phi$, and eigenvalues $\Lambda$ (sec. 4.1). Given two observations $\psi$ and $T\psi$ related by some unknown transformation $T$ (in this case, camera motion in a 3D scene), they are first encoded into latent functions $\mathcal{E}{}(\psi)$ and $\mathcal{E}{}(T\psi)$ and projected into the operator eigenbasis. An isometric functional map $\tau{}_{\Omega{}}$ is estimated between them, and used to map one to the other. Losses promote isometry-equivariance in the latent space, reconstruction of transformed latents, and distinct, low-multiplicity eigenvalues $\Lambda$, with the latter encouraging a diagonal as possible $\tau{}_{\Omega{}}$ (sec. 4.2).
@@ 4.2 Estimating $\tau$ and End-to-End Optimization
In experiments, we also consider a similar triplet regime as proposed in [34, 35], where the autoencoder is given triples of $T$-related observations $(\psi,T\psi,T^{2}\psi)$ (assuming $T$ is composable) and the estimated map $\tau$ between the encodings of $(\psi,T\psi)$ is used to form equivariance and reconstruction losses between $(T\psi,T^{2}\psi)$ and vice-versa. This works to prevent $\tau{}$ from “cheating” by encoding privileged information about the relationship between pairs beyond $T$, a property we show to be critical in enforcing a useful notion of latent equivariance. However, triples of $T$-related observations are rare in practical settings, and we show in experiments that a major benefit of our isometric regularization is that our multiplicity loss (promoting a diagonal-as-possible and thus sparse and compact $\tau{}_{\Omega{}}$) can serve as an effective substitute for access to triples.
@@
Figure 3: Approximating the Toric Laplacian from Shifted Images. Forced to map between shifted images (left), NIso regress an operator (center right) structurally similar to the toric Laplacian (right). Maps $\tau{}_{\Omega{}}{}$ between projected images are strongly diagonal (center left), with individual blocks (inset) preserving the subspaces spanned by eigenfunctions (center) sharing nearly the same eigenvalues.
4.3 A Simple Example: Approximating the Toric Laplacian
@@
We demonstrate that NIso are able to learn a compact representation of isometries that reflect the dynamics of transformations in world space. To do so, we consider pairs of observations formed by $16\times 16$ images from ImageNet [41] perturbed by random circular shifts along both spatial axes. Thus pairs are related by the isometries of the torus which commute with the Laplacian. Taking the encoder and decoder to be the identity map (making the equivariance and reconstruction losses equivalent) we optimize for $\mathbb{M}{},\Omega\in\mathbb{R}^{256\times 256}$ via the pairwise training procedure described in sec. 4.2. As seen in Fig 3, NIso regresses an operator with significant structural similarities to the Laplacian matrix formed by the standard $3\times 3$ stencil and an eigenspace that diagonalizes $\tau{}_{\Omega{}}{}$ between shifted images. As our approach is data-driven and our estimated maps are only approximately isometric, our learned operator and its eigenspectrum do not exactly correspond to the ground-truth Laplacian. Instead, we are able to characterize similar, non-trivial spatial relationships that are preserved under shifts.
@@ 5 Experiments
In this section, we provide empirical evidence through experiments that NIso 1) recovers a general-purpose latent space that can be exploited by isometry-equivariant networks to handle challenging symmetries (5.1, 5.2); and 2) NIso encodes information about transformations in world space through the construction of isometric maps in the latent space from which geometric quantities such as camera poses can be directly regressed (5.3). We provide reproducibility details in the supplement. We note that a consistent theme in our experiments are comparisons against the unsupervised variant of the NFT [35] (the semi-supervised variants cannot be applied because the symmetries we consider have no finite-dimensional IRs). Like our approach, the NFT seeks to relate latents via linear transformations though, it differs fundamentally in that maps are guaranteed additional structure only if the world space transformations are a compact group action. While not originally proposed by the authors, we evaluate it in place of our approach in the same self-supervised representation learning regimes discussed in sec. 4.4. Thus the role of these comparisons is to show that our proposed isometric regularization better and more consistently provides a tractable and informative latent space.
@@ 5.1 Homography-Perturbed MNIST
With this in mind, we perform three ablations. In the first, we train the NIso autoencoder in a triplet regime (4.2) made possible by the synthetic parameterization of $T$ as homographies. In the second and third, we train the autoencoder in the standard pairwise regime without considering the equivariance loss $\mathcal{L}_{E}$ and multiplicity loss $\mathcal{L}_{M}$, respectively. Additionally we compare the efficacy of our approach versus an NFT backbone pre-trained in the same manner. Last, we pre-train and evaluate a baseline backbone which considers only a reconstruction loss without respect to $T$-related pairs.
@@ 5.2 Conformal Shape Classification
Results are shown in Tab. 2, averaged over five randomly initialized pre-training and fine-tuning runs with standard errors. NIso outperforms the NFT and the autoencoder baseline in addition to Möbius Convolutions (MC) [7], a SoTA handcrafted spherical network equivariant to Möbius transformations. We consider this dataset to present a significant challenge as shape classes are roughly conformally-related and thus the maps between their spherical parameterizations are only approximated by Möbius transformations. This makes the performance of NIso particularly notable as it suggests that our framework has the potential to offer a more flexible and effective alternative than specialized, handcrafted equivariant networks whose underlying group-based architectures, though elegant, can only approximate inexact symmetries. Furthermore, while the NFT is relatively more competitive with the encoder backbone unfrozen, its performance and that of the baseline indicate that a lack of existing equivariant structure cannot easily be overcome in the fine-tuning phase.
@@ 5.3 Camera Pose Estimation from Real-World Video
[…]
We compare against the NFT using the same pre-training and fine-tuning procedure along with a transformer baseline. The latter is inspired by the recent success of such models in 3D vision tasks [46, 47], and uses the same ViT-based architecture proposed in DUSt3R [47], with the decoder modified to directly predict the parameters of the relative camera pose. As the weights of the NIso and NFT backbones are frozen during the fine-tuning phase, the transformer baseline is trained from scratch in the fine-turning phase to more directly compare the descriptiveness of the latent representations relative to the information contained in the raw observations.
@@
Results are shown in Tab. 3, averaged over five randomly initialized pre-training and fine-tuning runs with standard errors. All methods perform similarly with 00 frame skip but diverge afterwards, with NIso achieving significantly lower ATE values as the skip length increases. Notably, the ablative version of our method is consistently the worst performer, suggesting that isometric regularization is also critical to encode information about world space transformations. Overall the NFT achieves the second best performance, slightly outperforming NIso at 0 frame skip but diverging thereafter. However, the latent maps it recovers are neither sparse nor exhibit condensed structure (Tab.  3, center left), and we hypothesize that its inability to effectively regularize maps beyond linearity makes it difficult for the network to discover a consistent transformation model that applies across scales. This could cause the network to focus on a specific regime at the expense of others, which may explain its relatively strong performance at 0 frame skip. The transformer baseline is also highly flexible, and an analogous line of reasoning could explain its similar error profile.
@@ 6 Discussion
A key factor limiting the broader applicability of our approach to geometry processing and graph-based tasks is an inability to learn and transfer an operator between domains with varying connectivity. This could potentially be overcome by parameterizing $\Omega$ by a network over the simplices of a mesh or graph following [48], and we seek to address generalization between data representations in future work. Furthermore, we observe that the eigenvalues learned in our parameterization of $\Omega{}$ are not necessarily interpretable as frequencies in the classical sense as visibly smooth eigenfunctions may correspond to large eigenvalues, and those with fine details to small eigenvalues.
@@ Acknowledgements.
Use of the CO3Dv2 dataset is solely for benchmarking purposes and limited exclusively to the experiments described in sec. 5.3. We thank Ishaan Chandratreya for helpful discussions and David Charatan for his aesthetic oversight and for granting us permission to use his images captured for FlowMap [49] in Fig. 1-2.
@@ Appendix B Reproducibility
All code for replicating experiments will be made available at publication.
@@ B.1 Hardware
B.2 Approximating the Toric Laplacian
The model consists of a single NIso layer, which learns a full-rank operator and mass matrix, with the former parameterized by an orthogonal matrix of eigenvectors $\Phi\in\mathbb{R}^{256\times 256}$ and diagonal matrix of non-negative eigenvalues $\Lambda$. During training, input features are created by stacking 86 $16\times 16\times 3$ examples from ImageNet into a single $16\times 16\times 258$ tensor representing an observation $\psi$. Random circular shifts about both spatial axes are applied to form $T\psi$. As the encoder and decoders are the identity maps, the equivariance and reconstruction losses are equivalent, and the objective in Equation (11) is minimized with $\alpha=0$ and $\beta=0.5$. The model is trained for 100,000 iterations with a batch size of 1 using the AdamW optimizer [50] with a weight decay of $10^{-4}$. The learning rate follows a schedule consisting of a 2,000 step warm up from $0.0$ to $5\times 10^{-4}$ and afterwards decays to $5\times 10^{-5}$ via cosine annealing. Training takes approximately 6 hours.

### v2

@@ 1 Introduction
In this paper, we take a first step towards a class of models that learns to be equivariant to unknown and difficult geometric transformations in world space. We propose Neural Isometries (NIso), an autoencoder framework which learns to map the observation space to a latent space in which encodings are related by tractable, highly-structured linear maps whenever their corresponding observations are geometrically related in world space.
Specifically, observations are encoded into latents preserving their spatial dimensions. Images, for instance, can be encoded into latent functions defined over a lower resolution grid or the patch tokens of a ViT. For observations sharing some potentially unknown relationship in world space, we enforce their encodings be related by a functional map $\tau$ – a linear transformation on the space of latent functions. In particular, we require that $\tau{}$ is an isometry such that it preserves a learned inner product and commutes with a learned functional operator, in the sense that rigid body transformations commute with the Laplace operator in Euclidean space.
Neural Isometries exhibit unique properties that make them a promising step towards an architecture-agnostic regime for self-supervised equivariant representation learning. We experimentally validate two principle claims regarding the efficacy and applicability of our approach:
@@ 2 Related Work
Essential to our approach is the parameterization of transformations between observations in the latent space not as group representations, but instead as functional maps. Introduced in the seminal work of Ovsjanikov et al. [18], functional maps (FMs) provide a powerful medium for interpreting, manipulating, and constructing otherwise intractable mappings between 3D shapes via their realization as linear transformations on spaces of functions on meshes, forming the basis for state-of-the-art pipelines in shape matching and correspondence [19, 20, 21, 22, 23]. Beyond 3D shapes, FMs can be seen as a tool to parameterize transformations between observations viewed as functions, in the sense that images, for instance, are functions that map 2D pixel coordinates to RGB colors. Integral to their study and implementation is the Laplace-Beltrami operator (the generalization of the Laplace operator to function spaces on manifolds), and in particular the expression of FMs in its eigenbasis which can expose specific geometric properties of deformations. In particular, isometries (distance-preserving transformations) manifest as highly-structured matrices not unlike IRs, being orthogonal and commuting with the diagonal matrix of eigenvalues and are thus approximately block-diagonal. That said, FMs are recovered through regularized linear solves and as such lack the consistency inherent in the analytical expressions of IRs for compact groups. However, freed from nagging theoretical constraints, FMs have displayed remarkable representational capacity, well modeling a variety of highly-complex non-rigid deformations [24] including those without group structure, such as partial [25] and inter-genus correspondence [26]. Please see [27] for an outstanding introduction to functional maps.
@@
A number of recent approaches have proposed autoencoder frameworks wherein given symmetries in the base space manifest as simple operations in the latent space [28, 29, 30, 31, 32, 33, 34, 35]. Perhaps most similar to our approach is recent work on a Neural Fourier Transform (NFT) that has sought to manifest group actions in observation space as IRs in the the latent space  [36, 37]. These methods offer impressive theoretical guarantees under the conditions that the observed transformations are either known or are a group action, although these assumptions may not hold for real-world data with complex symmetries. In contrast, our method is wholly unsupervised, and assumes no knowledge of the transformations in observation space nor that they even form a group.
Figure 2: Overview of Neural Isometries (NIso). NIso learn a latent space where transformations of observations manifest as isometries, achieved by regularizing the functional maps $\tau$ between latents to commute with a learned operator $\Omega$, parameterized via its spectral decomposition into a mass matrix $\mathbb{M}$, eigenfunctions $\Phi$, and eigenvalues $\Lambda$ (sec. 4.1). Given two observations $\psi$ and $T\psi$ related by some unknown transformation $T$ (in this case, camera motion in a 3D scene), they are first encoded into latent functions $\mathcal{E}{}(\psi)$ and $\mathcal{E}{}(T\psi)$ and projected into the operator eigenbasis. An isometric functional map $\tau{}_{\Omega{}}$ is estimated between them, and used to map one to the other. Losses promote isometry-equivariance in the latent space, reconstruction of transformed latents, and distinct, low-multiplicity eigenvalues $\Lambda$, with the latter encouraging a diagonal as possible $\tau{}_{\Omega{}}$. An optional spectral dropout layer can be applied before the basis unprojection to encourage a physically meaningful ordering of the learned spectrum (sec. 4.2).
@@ 4.2 Estimating $\tau$ and End-to-End Optimization
In experiments, we also consider a similar triplet regime as proposed in [36, 37], where the autoencoder is given triples of $T$-related observations $(\psi,T\psi,T^{2}\psi)$ (assuming $T$ is composable) and the estimated map $\tau$ between the encodings of $(\psi,T\psi)$ is used to form equivariance and reconstruction losses between $(T\psi,T^{2}\psi)$ and vice-versa. This works to prevent $\tau{}$ from “cheating” by encoding privileged information about the relationship between pairs beyond $T$, a property we show to be critical in enforcing a useful notion of latent equivariance. However, triples of $T$-related observations are rare in practical settings, and we show in experiments that a major benefit of our isometric regularization is that our multiplicity loss (promoting a diagonal-as-possible and thus sparse and condensed $\tau{}_{\Omega{}}$) can serve as an effective substitute for access to triples.
Spectral Dropout.
While the composite loss in Equation (11) promotes latent equivariance and distinct eigenvalues, it does not, however, promote a physically meaningful ordering of the eigenvalues – i.e. that small eigenvalues correspond to smooth, low-frequency eigenfunctions and larger eigenvalues to eigenfunctions with sharper, high-frequency details. Though such an ordering amounts to a permutation in the spectral dimension and is not a necessary condition for the existence of diagonalized isometric maps, it could be potentially useful in downstream applications where a classical notion of eigenvalues as frequencies is desired.
To this end we propose an optional spectral dropout layer applied after computing the equivariance loss $\mathcal{L}_{E}$ and before the basis unprojection. The layer is implemented as follows: During training, there is a $50\%$ chance that dropout will be applied to given example in a batch. Then, for each such example, a spectral index $1<i\leq k$ is randomly chosen and all features with index $j\geq i$ are masked out. Intuitively, this forces the decoder to produce the best possible reconstructions with the lowest frequency eigenfunctions, which appear most often and thus must capture large scale features, while reserving higher frequencies for filling in fine details.
@@
Figure 3: Approximating the Laplacian. Forced to map between shifted images on the torus (first row, left) and rotated images on the sphere (second row, left), NIso regress operators (center right) structurally similar to the toric and spherical Laplacian (right). Maps $\tau{}_{\Omega{}}{}$ between projected images are strongly diagonal (center left), with individual blocks (inset) preserving the subspaces spanned by eigenfunctions (center, first $64$ shown) sharing nearly the same eigenvalues. These experiments result in the discovery of basis with the similar properties to the the toric and spherical harmonics. In particular, the estimated spherical $\tau_{\Omega}$ manifest exactly the same structure as the ground truth Wigner-D matrices corresponding to the rotation, with square blocks of size $(2\ell+1)\times(2\ell+1)$ for the $\ell$-th distinct eigenvalue. Please zoom in to view structural details.
4.3 A Simple Example: Approximating the Toric and Spherical Laplacians
@@
We demonstrate that NIso are able to learn a compact representation of isometries that reflect the dynamics of transformations in world space. To do so, we perform two experiments in which we consider pairs of observations formed by $16\times 16$ images from ImageNet [43] viewed functions on toric and spherical grids, and respectively transformed by random circular shifts and SO(3) rotations to form pairs. Thus pairs are related by the isometries of the torus and sphere which commute with the Laplacian on each domain. Taking the encoder and decoder to be the identity map (making the equivariance and reconstruction losses equivalent) we optimize for $\mathbb{M}{},\Omega\in\mathbb{R}^{256\times 256}$ via the pairwise training procedure described in sec. 4.2, including the use of spectral dropout. For the toric experiments, we learn a full-rank parameterization of $\Omega$ with $k=256$; for the sphere, we learn a low-rank approximation with $k=64$. As seen in Fig 3, NIso regresses operators with significant structural similarities to the Laplacian matrices formed by the standard $3\times 3$ stencil on the torus and the low rank approximation using the first $64$ spherical harmonics. In both cases, NIso recovers an eigenspace that diagonalizes $\tau{}_{\Omega{}}{}$ between shifted images, with eigenfunctions ordered by their energy. As our approach is data-driven and our estimated maps are only approximately isometric, our learned operator and its eigenspectrum do not perfectly correspond to the ground-truth Laplacian. Instead, we are able to characterize similar, non-trivial spatial relationships that are preserved under shifts.
@@ 5 Experiments
In this section, we provide empirical evidence through experiments that NIso 1) recovers a general-purpose latent space that can be exploited by isometry-equivariant networks to handle challenging symmetries (5.1, 5.2); and 2) NIso encodes information about transformations in world space through the construction of isometric maps in the latent space from which geometric quantities such as camera poses can be directly regressed (5.3). Here we pre-train NIso without spectral dropout, as the ordering of the learned spectrum is irrelevant in our target applications and we find it slightly decreases accuracy of the prediction head. We provide reproducibility details in the supplement in addition to experiments quantitatively evaluating the degree of learned equivariance. We note that a consistent theme in our experiments are comparisons against the unsupervised variant of the NFT [37] (the semi-supervised variants cannot be applied because the symmetries we consider have no finite-dimensional IRs). Like our approach, the NFT seeks to relate latents via linear transformations though, it differs fundamentally in that maps are guaranteed additional structure only if the world space transformations are a compact group action. While not originally proposed by the authors, we evaluate it in place of our approach in the same self-supervised representation learning regimes discussed in sec. 4.4. Thus the role of these comparisons is to show that our proposed isometric regularization better and more consistently provides a tractable and informative latent space.
@@ 5.1 Homography-Perturbed MNIST
With this in mind, we perform three ablations. In the first, we train the NIso autoencoder in a triplet regime (4.2) made possible by the synthetic parameterization of $T$ as homographies. In the second and third, we train the autoencoder in the standard pairwise regime without considering the equivariance loss $\mathcal{L}_{E}$ and multiplicity loss $\mathcal{L}_{M}$, respectively. Additionally we compare the efficacy of our approach versus an NFT backbone pre-trained in the same manner. Last, we pre-train and evaluate a baseline backbone which considers only a reconstruction loss without respect to $T$-related pairs, but trains with data-augmentation during the fine-tuning phase by applying randomly sampled homographies.
@@ 5.2 Conformal Shape Classification
Results are shown in Tab. 2, averaged over five randomly initialized pre-training and fine-tuning runs with standard errors. NIso outperforms the NFT and the autoencoder baseline (with random Möbius transformations applied during the fine-tuning phase) in addition to Möbius Convolutions (MC) [7], a SoTA handcrafted spherical network equivariant to Möbius transformations. We consider this dataset to present a significant challenge as shape classes are roughly conformally-related and thus the maps between their spherical parameterizations are only approximated by Möbius transformations. This makes the performance of NIso particularly notable as it suggests that our framework has the potential to offer a more flexible and effective alternative than specialized, handcrafted equivariant networks whose underlying group-based architectures, though elegant, can only approximate inexact symmetries. Furthermore, while the NFT is relatively more competitive with the encoder backbone unfrozen, its performance and that of the baseline indicate that a lack of existing equivariant structure cannot easily be overcome in the fine-tuning phase.
@@ 5.3 Camera Pose Estimation from Real-World Video
[…]
We compare against the NFT using the same pre-training and fine-tuning procedure along with a transformer baseline. The latter is inspired by the recent success of such models in 3D vision tasks [50, 51], and uses the same ViT-based architecture proposed in DUSt3R [51], with the decoder modified to directly predict the parameters of the relative camera pose. As the weights of the NIso and NFT backbones are frozen during the fine-tuning phase, the transformer baseline is trained from scratch in the fine-turning phase to more directly compare the descriptiveness of the latent representations relative to the information contained in the raw observations. We also compare against two strong representation learning baselines. We extract features from both images using two pre-trained state-of-the-art vision foundation models — DINOv2[47] and BeIT[48] — and pass the tokens into the modified DUSt3R-style decoder which is trained to predict the pose.
Results are shown in Tab. 3, averaged over five randomly initialized pre-training and fine-tuning runs with standard errors. All methods perform similarly with 00 frame skip but diverge afterwards, with NIso achieving significantly lower ATE values as the skip length increases. Notably, the ablative version of our method is consistently among the worst performers, suggesting that isometric regularization is also critical to encode information about world space transformations. Overall the NFT achieves the second best performance, slightly outperforming NIso at 0 frame skip but diverging thereafter. However, the latent maps it recovers are neither sparse nor exhibit condensed structure (Tab.  3, center left), and we hypothesize that its inability to effectively regularize maps beyond linearity makes it difficult for the network to discover a consistent transformation model that applies across scales. This could cause the network to focus on a specific regime at the expense of others, which may explain its relatively strong performance at 0 frame skip. The transformer baseline and representation learners are also highly flexible, and an analogous line of reasoning could explain their similar error profile.
@@ 6 Discussion
A key factor limiting the broader applicability of our approach to geometry processing and graph-based tasks is an inability to learn and transfer an operator between domains with varying connectivity. In addition, NIso does not explicitly handle partiality or occlusion between input pairs which are ubiquitous in real-world data and likely degrade its performance in the pose estimation task. We seek to address these limitations in future work.
@@ Acknowledgements.
Use of the CO3Dv2 dataset is solely for benchmarking purposes and limited exclusively to the experiments described in sec. 5.3. We thank Ishaan Chandratreya and Hyunwoo Ryu for helpful discussions and David Charatan for his aesthetic oversight and for granting us permission to use his images captured for FlowMap [52] in Fig. 1-2.
@@ Appendix B Reproducibility
Code and experiments are available at https://github.com/vsitzmann/neural-isometries.
@@ B.1 Hardware
B.2 Approximating the Laplacian
In the toric/spherical experiments, the model consists of a single NIso layer. which learns a full-rank/low-rank operator and mass matrix, with the former parameterized by an orthogonal matrix of eigenvectors $\Phi\in\mathbb{R}^{256\times 256}$/$\Phi\in\mathbb{R}^{256\times 64}$ and diagonal matrix of non-negative eigenvalues $\Lambda$. During training, input features are created by stacking 86 $16\times 16\times 3$ examples from ImageNet into a single $16\times 16\times 258$ tensor representing an observation $\psi$. In the toric experiments, random circular shifts about both spatial axes are applied to form $T\psi$. In the spherical experiments, the images are assumed to lie on a $16\times 16$ Driscoll-Healy spherical grid and random rotations in $\textrm{SO}(3)$ are sampled and applied to form $T\psi$. As the encoder and decoders are the identity maps, the equivariance and reconstruction losses are equivalent, and the objective in Equation (11) is minimized with $\alpha=0$ and $\beta=0.1$. The models is trained with spectral dropout for 100,000 iterations with a batch size of 1 using the AdamW optimizer [53] with a weight decay of $10^{-4}$. The learning rate follows a schedule consisting of a 2,000 step warm up from $0.0$ to $5\times 10^{-4}$ and afterwards decays to $5\times 10^{-5}$ via cosine annealing. Training takes approximately 6 hours.